# references.bib

### v1

@@ 3.2 Quantum Computing
Parallel to our work, Zappin et al. [zappin_when_2024] recently released a comprehensive study characterizing Hybrid Quantum-Classical issues discussed in Developer Forums, which represents the closest work to ours to date. However, they solely investigated discussions available on Xanadu Discussion Forums and on QC Stack Exchange. The issues they used do not originate from GitHub and do not follow the strict inclusion criteria we used as described in Section 4, hence there is no overlap between our two works. Their study gave an overview of current causes of issues in a software engineering perspective, gathering them around the following categories: Software Faults, Hardware/Simulator issues, Configuration issues, Developper Errors, Library and Plateform Issues. Our study characterises faults in Hybrid Quantum-Classical architectures around their structuring components, and identify their typical weaknesses, aiming to further understand their nature and stay relevant to future versions of such architectures. These two taxonomies are therefore complementary. They were conducted in parallel, without consulting one another, which highlights once more the relevance of such a work.
4 Methodology
Fig. 3 presents an overview of our methodology to construct the final taxonomy. We detail the methodology in the remainder of this section based on the steps depicted in Fig. 3.
Figure 3: Overview of the 6 steps that led to the construction of our Final Taxonomy.
@@ 1.1. Inclusion Criteria
We aim to comply with ACM guidelines for reproducibility [noauthor_artifact_nodate], which means a different team should be able to reproduce these faults using the same setup.
@@ 1.3. Designing a Search Query
We used GitHub search API to find Hybrid Quantum-Classical architecture faults and manually assess them. After an initial manual and manual search investigation, we discovered that NISQ algorithms are the only Hybrid Quantum-Classical architecture for which there are examples of public faults and bug fixes satisfying our criteria. We first searched ’NISQ’ and ’Noisy intermediate-Scale Quantum’. We found these keywords to be too generic. They mostly led to discussion around research papers, or theoretical questions about the nature of NISQ, with a vast majority of results not written in English. All the 718 results were directly excluded by at least one of our criteria, and often all four of them. Hence, we decided to search for specific algorithms designed for Hybrid Quantum-Classical architectures.
Most implementations require importing a package named after the algorithms. For example, in qiskit, we import a class named VQE to implement VQEs. Any source code including such an import, or any discussion mentioning the searched algorithms will be identified by this method. We, therefore, used the following keywords, which are the available algorithms according to a recent survey [bharti_noisy_2022]: VQE, VQA, Variational Quantum eigensolvers, Variational Quantum algorithm, Quantum Annealing, Gaussian boson sampling, Analog quantum simulation, Digital-analog quantum simulation and computation, Iterative quantum assisted eigensolver, Quantum Approximative Optimisation Algorithms, QAOA, Quantum Machine Learning, QML, Tensorflow Quantum, and TFQ.
@@
In this step, we performed the search and stored the results (2.1); applied our inclusion/exclusion criteria and performed first the manual filtration of the results by one author and then their labelling by two authors. Table 1 provides an overview of the number of issues from our initial search process (in the column marked Initial) and after each stage of cross-validation (columns marked "Filtration" and "Labelling").
@@ 2.2. Manual Filtration
We then excluded any issue that does not comply with one or more of our criteria: it does not concern a Hybrid Quantum-Classical architecture, it does not provide code or fix, or it is not in the English language. This process was performed manually by the first author for each of the 5072 bugs. As a result of applying the inclusion criteria, 529 faults remained in the dataset, as reported in column "Filtered" in Table 1.
The results of manual filtration show that ‘QML resulted in a high number of false positives. The reason behind this is that ‘QML’ stands not only for ’Quantum Machine Learning’, but also for ’QT Modeling Language’, which typically uses a file type ‘.qml’. Out of the obtained 2861 results, only 62 were related to Quantum programming, and only 6 were about Hybrid Quantum-Classical architectures. Pennylane, which is a common framework for quantum machine learning, is often imported using ‘import pennylane as qml’ as advised in their official documentation. Any program using this import without implementing a Hybrid Quantum-Classical architecture is caught by searching ‘QML’. Also, all remaining 6 issues were caught by other keywords. Our results also indicate that the abbreviations such as VQA instead of Variational Quantum Algorithms, lead to fewer false positives, as using the full names of algorithms typically leads to issues the majority of which lack any source code.
Table 1: Number of identified results after each stage: “Initial” refers to the initial number of results obtained through GitHub search, “Filtered” refers to the number of issues that satisfy the inclusion criteria, “Labelled” refers to the number of issues that were decided to be relevant during the manual labelling process by two authors.
@@ 3.2. Individual Labelling
Lastly, for each bug, we identify its symptom (such as crash, failure, slow performance, wrong output) and create a short label that summarises its root cause (for example, "suboptimal kernel building"). During the labelling procedure, we kept the list of already generated labels available to enable their reuse when necessary. We have labelled the 529 bugs in 8 rounds (targeting on average 60 bugs per round) and had a conflict resolution meeting after each round. We carefully tracked all conflicts (108 in total) and how they were resolved.
@@ 4.1. Selecting Interviewees
To ensure diversity in our panel of experts, especially between industry and academia, we selected and contacted experts working on Hybrid Quantum-Classical architectures at several leading quantum computing companies such as Quantum IBM, Quasar USA Quantum Blockchain Technologies, Google Quantum, and ClassicQ. We approached them through e-mail, as well as LinkedIn and ResearchGates private messaging systems. We received three positive answers and managed to secure another industry interview (a total of four) through personal contacts. One of the interviewees from the industry was identified by us due to their active participation in fixing several of the GitHub issues in our dataset and we contacted them on the LinkedIn platform. The remaining selected experts (seven in total) are active in academia. Two of them have a strong connection with the industry. Academic profiles vary from PhD student to Lecturer, both from Computer Science and Physics Departments with significant experience in Quantum Computing and especially in Hybrid Classical-Quantum architectures.
@@ 4.2. Designing the Interview Structure
We opted for a semi-structured interview to have a flexible and adaptive structure to elicit as much information from the experts as possible [seaman_qualitative_1999]. After establishing the scope of the study, and defining the key terms as advised in the pilot interviews, we opened with a broad question, namely, ’What kind of problems have you experienced developing Hybrid Quantum-Classical architectures?’. We tried to adapt the questions to the interviewees and re-direct the discussion to lead them to share the faults they have experienced. We prepared a script, available in the replication package, to ensure consistency in the content of each interview while being adaptive to elicit most information where the interviewee had more experience. When the interviewee mentioned any fault during the interview, we asked additional questions to identify all the characteristics required for our collection (in the library or not, symptoms, etc.)
@@ 5.1. Search Validation
We explored other quantum bug benchmarks [zhao_bugs4q_2021][zhao_empirical_2023] and selected all bugs that satisfied our inclusion criteria. We verified that all these bugs were included also in our dataset. Bugs4Q [zhao_bugs4q_2021] included only 3 bugs with a noise simulation, and none were Hybrid Quantum-Classical architecture faults. The second study [zhao_empirical_2023] focused on Quantum Machine Learning bugs. To validate our dataset, we checked all benchmark faults complying with our four criteria in the benchmark repositories and ensured our dataset included them. We selected 71 bugs in their repository, all of which were indeed already captured in our results.
@@ 5.1 Faults Dataset
Table 2: Overview of the final result of faults included in the dataset: “Level” is the fault location in the hybrid architecture; “Faults” is the total number of faults, differentiating betw. GitHub- and interview-sourced faults (GitHub+Interviews); “In Lib.” denotes the faults in the libraries and “Not in Lib.” are those in the code using the libraries (both differentiating Github- + Interview-sourced faults).
@@ 5.2 Final Taxonomy
Table 3: Faults happening at the parametrisation phase of the system, divided into 8 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
This category emerged from the interviews and accounts for 33% of the faults mentioned by experts as shown in Table 2. It consists of the conceptual mistakes made while translating the domain knowledge into an instance of a Hybrid Quantum-Classical architecture. This translation process is far from trivial and several faults that may be introduced in the translation process are listed in Table 4. In Section 2.3, we have introduced the key concepts used in this category, such as Hamiltonians, which is the representation of a quantum system and corresponds to a specific electronic structure. During the iteration process, this electronic structure corresponding to the represented system can be broken, leading to erroneous behaviours. More details can be found in Tilly et al.’s detailed review of VQE [tilly_variational_2022].
For instance, four experts mentioned Conceptual faults introduced by suboptimal Ansatz design as discussed in this interview. A poor Ansatz design can significantly affect the optimisation process: it will lead to performance issues and wrong outputs. It can be the root cause of Barren plateaus issues. Yet, finding an optimal Ansatz design is very challenging even for domain experts, it is problem-specific, and still an active area of research [tilly_variational_2022].
Figure 4: The final taxonomy, organised into 7 top-level categories in darker colours, divided into 47 sub-categories in lighter shades of the same colour.
Table 4: Faults happening at the conceptualisation phase, comprising 6 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
Table 5: Faults happening at the API level of the system, divided into 3 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
This category concerns the faults occurring in the classical code for optimisation. We initially named this category Training, refering to Humbatova et al.’s Taxonomy of Real Faults in Deep Learning Systems [humbatova_taxonomy_2019]. However, we realised at a later stage that Optimisation defines better this category in the context of Hybrid Quantum-Classical architectures. This category is still labelled as Training in the replication package to acknowledge the evolution of our study. It is the biggest category of our taxonomy, both in the numbers of faults and of sub-categories, accounting for about 25% of GitHub faults and 20% of interviews‘ faults, divided into 12 sub-categories, listed in Table 6. The large number and diversity of faults in this category are motivated by the corresponding complexity and size of the code: due to the noise in the current architectures, the amount of classical code for optimisation dominates the complexity and size of the quantum circuit. At each iteration, the measurements of the quantum circuit are stored as intermediate results, the cost function is updated, and the coefficients of each parametric gate and the weights of the Ansatz’s layers are updated for the next iterations. A more detailed explanation is available in Tilly et al.’s review of VQE [tilly_variational_2022].
@@
Table 6: Faults happening at the optimisation phase of the system, divided into 12 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
Table 7: Faults originating from the quantum circuit, divided into 8 sub-categories. Typical cause is the most used label in this category, and Nb is the number of faults from online resources and interviews, respectively.
@@
Table 8: Faults originating from the measurements, divided into 6 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
Table 9: Faults originating from the GPU integration, divided into 4 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
5.3 Validation Results
Table 10 displays the results of the validation survey. One of the seven participants answered "No" to the opening question ’Have you experienced Hybrid Quantum-Classical architecture faults’, so we decided to exclude their contribution as it did not fit the required expertise - they only had some high-level theoretical knowledge of Hybrid Quantum-Classical architectures. Not all participants filled every sub-questions about the severity and effort to solve per category; the answers received are summarised in Table 10. 4 participants declared having a Computer Science background, 2 came from Physics education, and one from Mathematics.
The results of the survey generally support the findings of the Github mining and the interviews: the identified categories of faults were experienced by at least 50% of the participants, with 4 categories experienced by 83% of them. All but one participant experienced at least half of the categories of faults. The aforementioned remaining participant has experienced 3 out of 7. When asked if they have encountered faults not present in the survey, two mentioned noise-induced errors - present as a sub-category of Measurements, and one mentioned a specific API problem that would fall into our API Misuse sub-category. The rest of the participants answered negatively. The survey provided additional insights that can be further used when developing testing, debugging, and repair techniques for these faults: critical faults seem to be concentrated in API, GPU, and measurement components. As we report below, our manual analysis also corroborates that the interfaces of components, particularly across the quantum and classical boundry are most prone to severe faults. The survey indicates that API faults are not only critical but difficult to resolve, while conceptualisation faults, which are mostly minor issues, are also very difficult to resolve due to the deep insight needed both in the domain and in the translation to quantum concepts. The following section details further our findings.
Table 10: Validation Survey Results. In each category, the “Answer” column indicates if the experts have encountered any such faults before. For those who have identified the faults in the category a further indication of the “Severity” (minor, major, or critical) of and the “Effort to Resolve” (low, medium, or high) the encountered fault are provided in the corresponding columns.

### v2

@@ 3.2. Quantum Computing
Parallel to our work, Zappin et al. (Zappin et al., 2024) recently released a comprehensive study characterizing Hybrid Quantum-Classical issues discussed in Developer Forums, which represents the closest work to ours to date. However, they solely investigated discussions available on Xanadu Discussion Forums and on QC Stack Exchange. The issues they used do not originate from GitHub and do not follow the strict inclusion criteria we used as described in Section 4, hence there is no overlap between our two works. Their study gave an overview of current causes of issues in a software engineering perspective, gathering them around the following categories: Software Faults, Hardware/Simulator issues, Configuration issues, Developper Errors, Library and Plateform Issues. Our study characterises faults in Hybrid Quantum-Classical architectures around their structuring components, and identify their typical weaknesses, aiming to further understand their nature and stay relevant to future versions of such architectures. These two taxonomies are therefore complementary. They were conducted in parallel, without consulting one another, which highlights once more the relevance of such a work.
3.3. Comparative Analysis of Our Study
Data Source
A key strength of our study lies in the diversity and rigor of its data sources. While most previous datasets either focus exclusively on synthetic benchmarks such as QASMBench (Luo et al., 2022; Li et al., 2023), on general-purpose quantum software like Bugs4Q (Zhao et al., 2021b) and QBugs (Campos and Souto, 2021), or on community-reported developer issues without validation such as the recent work by Zappin et al. (Zappin et al., 2024), our study draws on real-world Hybrid Quantum–Classical architectures faults mined directly from GitHub repositories verified and fixed by developers. We complemented this empirical evidence with curated expert knowledge obtained through targeted interviews and surveys. This multi-pronged data collection strategy allows our taxonomy to be grounded in both reproducible faults and practical insights from developers and researchers actively engaged in Hybrid Quantum-Classical architectures design and deployment.
Focus
Prior studies have focused on specific domains such as classical software engineering (e.g., Defects4J (Just et al., 2014)), deep learning systems (e.g., Humbatova et al. (Humbatova et al., 2019); Thomas et al. (Thomas et al., 2025)), or specialised subfields such as Quantum Machine Learning (e.g., Zhao et al. (Zhao et al., 2023)). In contrast, our work targets the emerging domain of Hybrid Quantum–Classical architectures, where quantum and classical components interact within the same computational workflow. This focus reveals new categories of faults that cannot be adequately captured by existing taxonomies developed for purely classical or purely quantum contexts. By centring on real Hybrid architectures, our taxonomy captures the distinctive fault patterns and interaction failures characteristic of hybrid systems, providing researchers and practitioners with a timely and practically relevant foundation for advancing testing and debugging methodologies in Hybrid architectures engineering.
Validation
A defining strength of our work lies in its rigorous validation methodology. Whereas prior studies have typically relied on manual curation (e.g., Bugs4Q (Zhao et al., 2021b); QBugs (Campos and Souto, 2021)) or literature-based triangulation (e.g., Gill et al. (Gill et al., 2022)), our approach introduces a more systematic, multi-stage validation pipeline. This pipeline combines cross-validation with existing datasets, triangulation across heterogeneous sources—including GitHub issue reports, developer interviews, and targeted surveys—and expert assessment obtained through both interviews and a dedicated round of validation surveys. This methodological depth ensures that our taxonomy is empirically grounded and verified for accuracy and practical relevance by a diverse set of practitioners and researchers across multiple quantum-computing subfields. To reduce false positives (i.e., issues that do not correspond to actual faults), we restricted our dataset to closed issues, guaranteeing that each entry reflects a confirmed and resolved fault. This choice enhances the reliability and reproducibility of our results and aligns with best practices adopted in previous peer-reviewed software-fault datasets (Zhao et al., 2021 (Zhao et al., 2021b); Campos et al., 2021 (Campos and Souto, 2021); Luo et al., 2022 (Luo et al., 2022); Li et al., 2023 (Li et al., 2023); Zappin et al., 2024 (Zappin et al., 2024)).
Limitation
Despite these strengths, we acknowledge that our taxonomy, while broad and empirically grounded, is inherently shaped by the current state of hybrid quantum-classical technologies and the faults that have been reported and documented to date. As quantum hardware, software frameworks, and hybrid orchestration techniques evolve, new classes of faults will inevitably emerge. This limitation underscores the importance of designing our taxonomy to be extensible and adaptable to future architectures — a design choice that we have explicitly prioritised to ensure its long-term relevance and usefulness to the research and development community.
Table 1. Summary of Related Work and Positioning of Our Study
Study / Dataset Data Source Focus Validation Steps Gap / Limitation Defects4J (Just et al., 2014) Java GitHub Repositories Real faults dataset for testing and debugging Extensive dataset curation and cross-validation with multiple Java projects Not applicable to quantum or Hybrid architectures Humbatova et al. (Humbatova et al., 2019), Thomas et al. (Thomas et al., 2025) GitHub, Stack Overflow Taxonomy of real faults in deep learning applications Expert-based evaluation of taxonomy; empirical validation on DL repositories Does not cover quantum or Hybrid software architectures Zhao et al. (Bugs4Q) (Zhao et al., 2021b) GitHub (Qiskit) Benchmark of bugs in quantum programs Manual curation of issues and patches from Qiskit repositories Contains only a single hybrid-related bug; not tailored to Hybrid architectures Campos et al. (QBugs) (Campos and Souto, 2021) GitHub (Qiskit, etc.) Catalog of quantum software bugs Validation through comparison with prior datasets and literature review Focuses on general quantum software; lacks hybrid specificity Zappin et al. (Zappin et al., 2024) Developer Forums (e.g., Xanadu, StackExchange) Developer-reported issues and discussions Descriptive qualitative validation via discussion analysis; no systematic cross-validation Lacks GitHub-based empirical faults; no validation or fault taxonomy Gill et al. (Gill et al., 2022) Literature Review Research priorities in quantum software engineering Literature triangulation and expert interpretation Mentions Hybrid architectures, but lacks empirical fault analysis and testing focus Zhao et al. (Zhao et al., 2023) GitHub (QML Frameworks) Empirical analysis of quantum machine learning software faults Manual inspection of commits and issue reports; triangulation with literature Concentrated on QML; not representative of general hybrid algorithms Miranskyy et al., Paltenghi et al., Metwalli et al., Pontolillo et al. (Miranskyy and Zhang, 2019; Paltenghi and Pradel, 2022; Metwalli and Van Meter, 2022; Pontolillo and Mousavi, 2023) Various (Literature, Tools) Highlight challenges in quantum software testing Conceptual validation through expert discussions and case studies in prior literature Do not address real-world faults or hybrid-specific taxonomies Luo et al., Li et al. (QASMBench) (Luo et al., 2022; Li et al., 2023) GitHub, Benchmarks Dataset of quantum bugs and low-level NISQ benchmark suite Benchmark validation against standard quantum workloads and reproducibility studies Focus on bug benchmarking and circuit-level simulation; not hybrid-focused Our Study GitHub + Manual Curation + Interviews and Survey Taxonomy of real faults in Hybrid architectures Cross-validation, validation with other datasets, expert interviews, and expert survey-based validation First empirical taxonomy based on real hybrid faults; grounded in current implementations and extensible to future architectures
Conclusion
In summary, compared to prior datasets and taxonomies (as outlined in Table 1), our study distinguishes itself through its diverse and reliable data sources, its focused attention on hybrid architectures, and its rigorous validation process that integrates expert perspectives with empirical evidence. By addressing the current gap in hybrid fault taxonomies and providing a reproducible dataset enriched by expert insight, our work lays a foundational basis for advancing testing strategies and improving the robustness of Hybrid quantum-classical software systems. We further detail the practical applications and future directions of our taxonomy in Section 6.5.
4. Methodology
Fig. 4 presents an overview of our methodology to construct the final taxonomy. To ensure the scientific rigour of our process, we incorporated several validation steps. These included a comprehensive cross-validation of our fault dataset by the two independent authors; expert interviews to diversify the origins of our dataset, with transcripts analysed separately and then discussed jointly by the authors; and a separate survey with an independent group of experts, used solely as a final validation layer. The two authors involved in the cross-validation are a P.h.D. student specialising in Quantum Computing and a Lecturer with expertise in Machine Learning and Software Engineering. They came from different institutions and have distinct backgrounds, ensuring their independence. Each validation step was carefully documented and is fully accessible in the replication package. We detail the methodology in the remainder of this section based on the steps depicted in Fig. 4.
Figure 4. Overview of the 6 steps that led to the construction of our Final Taxonomy.
@@ 1.1. Inclusion Criteria
We aim to comply with ACM guidelines for reproducibility (1), which means a different team should be able to reproduce these faults using the same setup.
@@ 1.3. Designing a Search Query
We used GitHub search API to find Hybrid Quantum-Classical architecture faults and manually assess them. After an initial manual search investigation, we discovered that NISQ algorithms are the only Hybrid Quantum-Classical architecture for which there are examples of public faults and bug fixes satisfying our criteria. We first searched ’NISQ’ and ’Noisy intermediate-Scale Quantum’. We found these keywords to be too generic. They mostly led to discussion around research papers, or theoretical questions about the nature of NISQ, with a vast majority of results not written in English. All the 718 results were directly excluded by at least one of our criteria, and often all four of them. Hence, we decided to search for specific algorithms designed for Hybrid Quantum-Classical architectures.
Most implementations require importing a package named after the algorithms. For example, in qiskit, we import a class named VQE to implement VQEs. Any source code including such an import, or any discussion mentioning the searched algorithms will be identified by this method. We, therefore, used the following keywords, which are the available algorithms according to a recent survey (Bharti et al., 2022): VQE, VQA, Variational Quantum eigensolvers, Variational Quantum algorithm, Quantum Annealing, Gaussian boson sampling, Analog quantum simulation, Digital-analog quantum simulation and computation, Iterative quantum assisted eigensolver, Quantum Approximative Optimisation Algorithms, QAOA, Quantum Machine Learning, QML, Tensorflow Quantum, and TFQ.
@@
In this step, we performed the search and stored the results (2.1); applied our inclusion/exclusion criteria and performed first the manual filtration of the results by one author and then their labelling by two authors. Table 2 provides an overview of the number of issues from our initial search process (in the column marked Initial) and after each stage of cross-validation (columns marked ”Filtration” and ”Labelling”).
@@ 2.2. Manual Filtration
We then excluded any issue that does not comply with one or more of our criteria: it does not concern a Hybrid Quantum-Classical architecture, it does not provide code or fix, or it is not in the English language. The manual filtration process was carried out by the first author, who opened and examined each of the 5072 issues extracted by the script individually. For each issue, the author read the title, body, and available code snippets (if any), and excluded any issue that clearly violated one or more of the inclusion criteria—such as being unrelated to Hybrid Quantum-Classical programming, lacking technical content, or being written in a non-English language. This step aimed to discard only obviously irrelevant issues, so as to preserve sensitivity. In cases where there was any ambiguity or doubt regarding the relevance of an issue, it was retained for further discussion and validation during the cross-review process between both authors in the next stage. As a result of applying the inclusion criteria, 529 faults remained in the dataset, as reported in column ”Filtered” in Table 2.
We note that this filtration followed a conservative benchmarking principle: exclude only those issues that were clearly non-qualifying, and defer ambiguous cases to the second-stage validation. The results of manual filtration show that ‘QML’ resulted in a high number of false positives. The reason behind this is that ‘QML’ stands not only for ’Quantum Machine Learning’, but also for ’QT Modeling Language’, which typically uses a file type ‘.qml’. Out of the obtained 2861 results, only 62 were related to quantum programming, and only 6 were about Hybrid Quantum-Classical architectures. Pennylane, which is a common framework for quantum machine learning, is often imported using ‘import pennylane as qml’ as advised in their official documentation. Any program using this import without implementing a Hybrid Quantum-Classical architecture is caught by searching ‘QML’. Also, all remaining 6 issues were caught by other keywords.
Our results also indicate that the abbreviations such as VQA instead of Variational Quantum Algorithms lead to fewer false positives, as using the full names of algorithms typically leads to issues the majority of which lack any source code.
Table 2. Number of identified results after each stage: “Initial” refers to the initial number of results obtained through GitHub search, “Filtered” refers to the number of issues that satisfy the inclusion criteria, “Labelled” refers to the number of issues that were decided to be relevant during the manual labelling process by two authors.
@@
Independence of the two authors
To mitigate potential bias arising from shared research affiliation, we implemented several measures to ensure independence and objectivity in the fault classification process. The two authors involved came from distinct disciplinary backgrounds (software engineering and quantum computing), had no prior working relationship at the time of the study, and conducted independent reviews and labelling of fault instances before resolving discrepancies through structured discussion. The classification task was inherently technical and systematic, with predefined labels and coding schemes, minimizing subjective interpretation. Furthermore, the resulting taxonomy underwent external validation through expert interviews and a follow-up survey, providing additional layers of triangulation and verification.
@@ 3.2. Individual Labelling
Lastly, for each bug, we identify its symptom (such as crash, failure, slow performance, wrong output) and create a short label that summarises its root cause (for example, ”suboptimal kernel building”). During the labelling procedure, we kept the list of already generated labels available to enable their reuse when necessary. We have labelled the 529 bugs in 8 rounds (targeting on average 60 bugs per round) and had a conflict resolution meeting after each round. We carefully tracked all conflicts (108 in total) and how they were resolved.
@@ 4.1. Selecting Interviewees
To ensure diversity in our panel of experts, especially between industry and academia, we selected and contacted experts working on Hybrid Quantum-Classical architectures at several leading quantum computing companies such as Quantum IBM, Quasar USA Quantum Blockchain Technologies, Google Quantum, and ClassicQ. We approached them through e-mail, as well as LinkedIn and ResearchGate’s private messaging systems. We received three positive responses and managed to secure another industry interview (a total of four) through personal contacts. One of the interviewees from the industry was identified by us due to their active participation in fixing several of the GitHub issues in our dataset, and we contacted them on the LinkedIn platform.The remaining selected experts (seven in total) are active in academia. Two of them have a strong connection with the industry. Academic profiles vary from PhD student to Professor, both from Computer Science and Physics Departments, with significant experience in Quantum Computing and especially in Hybrid Quantum-Classical architectures.To ensure a structured analysis, all experts were explicitly categorised based on their academic vs. industrial affiliation and disciplinary background. We directly asked participants about their expertise, current role, and experience during the interviews, as can be verified in the transcripts provided in our replication package (see the “Interview” folder). This categorisation allowed us to contextualise their feedback appropriately and ensured that insights from both academic and industrial perspectives were incorporated into the taxonomy refinement process.
@@ 4.2. Designing the Interview Structure
We opted for a semi-structured interview to have a flexible and adaptive structure to elicit as much information from the experts as possible (Seaman, 1999). After establishing the scope of the study, and defining the key terms as advised in the pilot interviews, we opened with a broad question, namely, ’What kind of problems have you experienced developing Hybrid Quantum-Classical architectures?’. We tried to adapt the questions to the interviewees and re-direct the discussion to lead them to share the faults they have experienced. We prepared a script, available in the replication package, to ensure consistency in the content of each interview while being adaptive to elicit most information where the interviewee had more experience. When the interviewee mentioned any fault during the interview, we asked additional questions to identify all the characteristics required for our collection (in the library or not, symptoms, etc.)
@@ 5.1. Search Validation
We explored other quantum bug benchmarks (Zhao et al., 2021b)(Zhao et al., 2023) and selected all bugs that satisfied our inclusion criteria. Given that those benchmarks include fixed bugs, they all comply with criteria 2 to 4, which means we selected all bugs that occurred in Hybrid Quantum Classical architectures. We verified that all these bugs were included also in our dataset. Bugs4Q (Zhao et al., 2021b) included only 3 bugs with a noise simulation, and none were Hybrid Quantum-Classical architecture faults. The second study (Zhao et al., 2023) focused on Quantum Machine Learning bugs. To validate our dataset, we checked all benchmark faults complying with our four criteria in the benchmark repositories and ensured our dataset included them. We selected 71 bugs in their repository, all of which were indeed already captured in our results.
@@ 5.1. Faults Dataset
Table 3. Overview of the final result of faults included in the dataset: “Level” is the fault location in the Hybrid architecture; “Faults” is the total number of faults, differentiating betw. GitHub- and interview-sourced faults (GitHub+Interviews); “In Lib.” denotes the faults in the libraries and “Not in Lib.” are those in the code using the libraries (both differentiating Github- + Interview-sourced faults).
@@ 5.2. Final Taxonomy
Table 4. Faults happening at the parametrisation phase of the system, divided into 8 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
This category emerged from the interviews and accounts for 33% of the faults mentioned by experts as shown in Table 3. It consists of the conceptual mistakes made while translating the domain knowledge into an instance of a Hybrid Quantum-Classical architecture. This translation process is far from trivial and several faults that may be introduced in the translation process are listed in Table 5. In Section 2.3, we have introduced the key concepts used in this category, such as Hamiltonians, which is the representation of a quantum system and corresponds to a specific electronic structure. During the iteration process, this electronic structure corresponding to the represented system can be broken, leading to erroneous behaviours. More details can be found in Tilly et al.’s detailed review of VQE (Tilly et al., 2022).
For instance, four experts mentioned Conceptual faults introduced by suboptimal Ansatz design as discussed in Interview 5. A poor Ansatz design can significantly affect the optimisation process: it will lead to performance issues and wrong outputs. It can be the root cause of Barren plateaus issues. Yet, finding an optimal Ansatz design is very challenging even for domain experts, it is problem-specific, and still an active area of research (Tilly et al., 2022).
Table 5. Faults happening at the conceptualisation phase, comprising 6 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
Figure 5. The final taxonomy, organised into 7 top-level categories in darker colours, divided into 47 sub-categories in lighter shades of the same colour.
@@
Table 6. Faults happening at the API level of the system, divided into 3 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
This category concerns the faults occurring in the classical code for optimisation. We initially named this category Training, refering to Humbatova et al.’s Taxonomy of Real Faults in Deep Learning Systems (Humbatova et al., 2019). However, we realised at a later stage that Optimisation defines better this category in the context of Hybrid Quantum-Classical architectures. This category is still labelled as Training in the replication package to acknowledge the evolution of our study. It is the biggest category of our taxonomy, both in the numbers of faults and of sub-categories, accounting for about 25% of GitHub faults and 20% of interviews‘ faults, divided into 12 sub-categories, listed in Table 7. The large number and diversity of faults in this category are motivated by the corresponding complexity and size of the code: due to the noise in the current architectures, the amount of classical code for optimisation dominates the complexity and size of the quantum circuit. At each iteration, the measurements of the quantum circuit are stored as intermediate results, the cost function is updated, and the coefficients of each parametric gate and the weights of the Ansatz’s layers are updated for the next iterations. A more detailed explanation is available in Tilly et al.’s review of VQE (Tilly et al., 2022).
@@
Table 7. Faults happening at the optimisation phase of the system, divided into 12 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
Table 8. Faults originating from the quantum circuit, divided into 8 sub-categories. Typical cause is the most used label in this category, and Nb is the number of faults from online resources and interviews, respectively.
@@
Table 9. Faults originating from the measurements, divided into 6 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
Table 10. Faults originating from the GPU integration, divided into 4 sub-categories. Typical cause is the most used label in this category, Nb is the number of faults from online resources and interviews respectively.
@@
5.3. Validation Results
Table 11 displays the results of the validation survey. One of the seven participants answered ”No” to the opening question ’Have you experienced Hybrid Quantum-Classical architecture faults’, so we decided to exclude their contribution as it did not fit the required expertise - they only had some high-level theoretical knowledge of Hybrid Quantum-Classical architectures. Not all participants filled every sub-question about the severity and effort to solve per category; the answers received are summarised in Table 11. 4 participants declared having a Computer Science background, 2 came from Physics education, and one from Mathematics.
The results of the survey generally support the findings of the Github mining and the interviews: the identified categories of faults were experienced by at least 50% of the participants, with 4 categories experienced by 83% of them. All but one participant experienced at least half of the categories of faults. The aforementioned remaining participant has experienced 3 out of 7. When asked if they have encountered faults not present in the survey, two mentioned noise-induced errors - present as a sub-category of Measurements, and one mentioned a specific API problem that would fall into our API Misuse sub-category. The rest of the participants answered negatively. The survey provided additional insights that can be further used when developing testing, debugging, and repair techniques for these faults: critical faults seem to be concentrated in API, GPU, and measurement components. As we report below, our manual analysis also corroborates that the interfaces of components, particularly across the quantum and classical boundary are most prone to severe faults. The survey indicates that API faults are not only critical but difficult to resolve, while conceptualisation faults, which are mostly minor issues, are also very difficult to resolve due to the deep insight needed both in the domain and in the translation to quantum concepts. We acknowledge that the relatively small sample size of the survey limits its statistical power and generalisability. Recruiting experts in the highly specialised field of Hybrid Quantum-Classical computing remains a significant challenge, especially as survey participants were independent from the interviewees. Therefore, the survey is intended primarily as a qualitative validation to confirm consistency with findings from empirical analysis and expert interviews, rather than to support formal statistical conclusions. The following section details further our findings.
Table 11. Validation Survey Results. In each category, the “Answer” column indicates if the experts have encountered any such faults before. For those who have identified the faults in the category a further indication of the “Severity” (minor, major, or critical) of and the “Effort to Resolve” (low, medium, or high) the encountered fault are provided in the corresponding columns.